4. Find the area of a triangle two silks of which are 18 cm and 10 cm and the perimeter is 42 cm.

Grade:12th pass

2 Answers

Pawan Prajapati
3 years ago
Peri.met.er = 42 cm. Third side = (42 -18- 10) cm = 14 cm. 42 s = 2 = 21 cm ... Area = 21(21- 18X21-10)(21-14) cm2 = 21x3x llx 7 cm2 = 21.Jii cm2•
ushan pathak
2 years ago
Peri.met.er = 42 cm. Third side = (42 -18- 10) cm = 14 cm. 42 s = 2 = 21 cm ... Area = 21(21- 18X21-10)(21-14) cm2 = 21x3x llx 7 cm2 = 21.Jii cm2•